OpenStudy (anonymous):

Find k(3) for the function k(n) = 4n2. A 36 B 24 C 10 D 49

jimthompson5910 (jim_thompson5910):

replace each copy of 'n' with 3 to get k(n) = 4n^2 k(3) = 4(3)^2 k(3) = ???